# Free3D: Consistent Novel View Synthesis without 3D Representation

### v1

@@ 3.2 View Consistent Rendering
Given a source image $x^{\text{src}}$, our goal is to render a series of consistent novel views $x^{i}$. While the camera encoding technique of Sec. 3.1 significantly enhances the pose accuracy of our model, if images are sampled independently, they will almost never be visually consistent due to the intrinsic ambiguity of the reconstruction task. This is visible as temporal flickering in the rendered 3D video (as shown in supplemental videos). To remove or at least greatly mitigate this problem, we propose to sample images jointly as well as to share noise among them.
@@
We first adapt the frame attention, a well-established method in video diffusion models [64, 30, 8, 26, 80], to capture temporal dependencies across different views. As shown in Fig. 2(c), given a 5D latent $z\in\mathbb{R}^{B\times v\times c\times h\times w}$, we initially reshape it to $z\in\mathbb{R}^{(B\times h\times w)\times v\times c}$, resulting in $batch\times height\times width$ sequences at the length of $views$. Subsequently, this reshaped latent is passed through the pseudo-3D attention module to calculate the similarity across different views. Since this attention layer operates across views but separately for each spatial location, the computational and memory costs are quite low (Tab. 1). Similar to ray conditioning, we inject multi-view attention at each level of the UNet $\hat{\epsilon}_{\theta}$.
@@
At test time, views are sampled using the backward process Eq. A.2 starting from a normally-distributed, yet randomly-sampled noise vector $x_{T}$. The idea is that the noise distribution is a representation of the data distribution, and models aleatoric uncertainty. In our case, while we are interested in generating multiple and diverse solutions, we still wish the different views to be internally consistent for each $360^{\circ}$ rendered video.
In order to reduce the variance between different views, we propose to start sampling each view from the same noise vector $x_{T}$. The network $\epsilon_{\theta}$ still generates different views because it is conditioned on the camera parameters. It can also generate different reconstructions by resampling $x_{T}$. However, the variance between views is thus reduced. Noise sharing can be justified by noting that the network $\hat{\epsilon}_{\theta}(z_{t},t,y)$ is a continuous function of both $z_{t}$ and $y$ [79]. Besides, different from [72], we do not specially set the noise schedule in different time steps $t$. More complex designs and training strategies have the potential to improve performance but are not the focus of this work.
@@ 3.3 Learning formulation
where each training sample $(Z_{0},z^{\text{src}},\mathbf{P})$ consist of $N$ encoded target views $Z_{0}=\{\mathcal{E}(x^{i})\}_{i=1}^{N}$, the encoded source view $z^{\text{src}}=\mathcal{E}(x^{\text{src}})$ and the viewpoints $\mathbf{P}$. The network is conditioned on the source view and cameras, i.e., $y=(z^{\text{src}},\mathbf{P})$ and, as in Fig. 2, estimates the noise for all target views jointly. This information is incorporated as discussed above via ray conditioning modulation; following [43], we also concatenate the input image code $z^{\text{src}}$ to $z_{t}$ along the channel dimension and add the CLIP [55] encoding.
Figure 3: Perceptual Path Length Consistency (PPLC). To partly compensate for the viewpoint change, the second image is rectified w.r.t. the first before comparison. To illustrate the importance of using rectification, the figure shows two objects in a large azimuth $\phi:57.6^{\circ}$. The top row shows to the left an ideally-rendered image pair, which however attains a large LPIPS loss due to the viewpoint change. To the right, rectification reduces this score. The bottom row shows the opposite, where a pair of incorrectly rendered views has its LPIPS loss increased by rectification.
@@ 3.4 Perceptual Path Length Consistency
To quantify the consistency across different novel views, recent methods [76, 43, 77, 83] trained NeRF models from the sampled views and evaluated them on the remaining views. However, this requires training a NeRF for each test instance, which is unfeasible for large-scale testing required for proper open-set NVS evaluation. Here, we propose to use instead the pairwise perceptual distance [88] to measure the consistency between generated views. In particular, we first subdivide the ${}^{\circ}$ rendering path into linear segments and then calculate the LPIPS score between two neighboring generated images $x^{i}$ and $x^{i+1}$. Naturally, these images differ due to the different viewpoints (Fig. 3 to row). We partly compensate for the viewpoint change by rectifying [27] the second image w.r.t. the first. We then define Perceptual Path Length Consistency (PPLC) of one rendered sequence $\{x^{i}\}_{i=1}^{N}$ as follows:
@@ 4 Experiments
Table 1: Comparison with SoTA methods on all 7,729 objects in Objaverse test-set. Recent works like [76, 43, 17, 77] were originally evaluated using a subset of this data only due to the cost of training additional 3D models. Unlike them, we directly evaluate the model on whole test-set without using additional 3D networks. The inference time is for rendering a single target view and a $360^{\circ}$ video, respectively.
We compare Free3D to state-of-the-art open-ended single-image NVS methods and also ablate our design choices.
@@ 4.1 Experimental Details
We follow [43, 77, 44] and assess the NVS quality by comparing the generated images and the ground-truth views at different levels of granularity, including pixel-level Peak Signal-to-Noise Ration (PSNR), patch-level Structure SIMilarity index (SSIM) and feature-level Learned Perceptual Image Patch Similarity (LPIPS) [88] [^1] https://github.com/richzhang/PerceptualSimilaritysqueeze”-net., and dataset-level Fréchet Inception Distance (FID) [28] [^2] https://github.com/GaParmar/clean-fidclip_vit_b_32”-net.. Some of these metrics look at the statistics of individually-reconstructed images rather than exact reconstructions, which is the correct approach given that the reconstruction task is inherently ambiguous; however, for the same reason, they are not suitable for assessing multi-view consistency. Instead, we use PPLC (Sec. 3.4) to evaluate consistency. To do so, we generate 50-frame videos along a pre-defined circular camera trajectory for each object and then compute the PPLC score between neighbouring frames.
@@
We compare Free3D to the state-of-the-art Zero-1-to-3 [43]${}_{\text{ICCV'2023}}$ and the follow-ups Zero123-XL [17]${}_{\text{NeurIPS'2023}}$, SyncDreamer [44]${}_{\text{arXiv'2023}}$ and Consistent123 [77]${}_{\text{arXiv'2023}}$, which are built on Zero-1-to-3. While noticed other concurrent works in arXiv for NVS [84, 35, 83], but, as no codes are available yet, we could not re-implement and re-train all of them for comparison.
@@ 4.2 Assessing Quality
Table 3: Ablations of Free3D design choices. Here, to reduce the computational cost and testing time, we only evaluate 2,000 instances in Objaverse dataset, instead of running the whole dataset with 7,729 for 50-frame $360^{\circ}$ video rendering. Therefore, the PPLC score is slight different from the values reported in Tab. 1.
@@
Qualitative results on various categories are visualized in Fig. 4. Free3D achieves better results even under challenging viewpoints with very different categories. SyncDreamer [44] uses an explicit volumetric representation representation, but can only generate views with fixed elevation ($30^{\circ}$), leading to worse results on synthesizing arbitrary target views. While the concurrent Consistent123 [77] aims to improve rendering consistency with a version of multi-view attention, they cannot directly improve pose accuracy. The Zero123-XL [17] learns to capture pose better than the baseline [43] by training on the larger Objaverse-XL, but the pose is still not very accurate. Because of the RCN layer, Free3D shows no such pose errors and results in better images.
@@ 4.3 Assessing Generalization
In Tab. 2, Free3D, which uses the same training set as Zero-1-to-3 [43], outperforms the baseline and all state-of-the-art variants, including Zero123-XL [17], which is trained on a larger 3D dataset, and SyncDreamer [44], which utilizes a heavier 3D volumetric representation. SyncDreamer improves the baseline on OmniObject3D, which has small elevation views change, while achieving worse results on GSO with random elevation angles. The concurrent Consistent123 [77] is also trained multiple views jointly, with additional cross-view attention. However, their relative improvement is limited. Table 2 shows that Free3D achieves very substantial improvements on all instantiations on both OmniObject3D and GSO datasets. This further indicates that ray conditioning can successfully improve the pose accuracy of the target view.
@@
A qualitative comparison is given in Fig. 5 (more in supplemental Figs. C.3 and C.4). Though Free3D is only trained on objaverse dataset, it works quite well in the open-set setting. Moreover, it shows significantly better results than all state-of-the-art models, and even better than the Zero123-XL [17], a concurrent model trained on much larger datasets.
@@ 4.4 Assessing 3D consistency
To assess 3D consistency, we render a $360^{\circ}$ video with the fixed elevation angle $\theta:=0^{\circ}$ and $50$ azimuth angles uniformly sampled in $[0^{\circ},360^{\circ}]$. The quantitative results are reported in Tabs. 1 and 2 using the proposed PPLC score, and the qualitative results are shown in Figs. 1 and 6. Consistent123 [77] still requires training additional NeRF to achieve 3D consistent video, while the directly rendered videos are still flickering. SyncDreamer [44] obviously improves view consistency by using an explicit 3D volume representation. However, it is trained only on a fixed elevation angle and can generate only 16 fixed frames for a video in its code. Interestingly, Free3D, by using multi-view attention and multi-view noise sharing is nearly as effective (while being much cheaper). This is also partly due to ray conditioning, which can capture more precise poses of the target view and thus reduce ambiguity. The additional results are provided as videos in the supplemental materials.
@@ 4.5 Ablation Study
We run a number of ablations to analyse Free3D, testing also the variants described in Sec. 3. Results are shown in Tab. 3 and discussed in detail next.
@@ Acknowledgements.
We use the Objaverse [18], OmniObject3D [81], and Google Scanned Object datasets (GSO) [20] following their terms and conditions. These datasets contain synthetic or scanned 3D objects, but, as far as we could determine, no personal data. For further details on ethics, data protection, and copyright please see https://www.robots.ox.ac.uk/~vedaldi/research/union/ethics.html.
@@ Appendix B Experiment Details
${}^{\circ}$ Video Rendering.
To render a ${}^{\circ}$ video, we establish a circle trajectory by uniformly subdividing the azimuth $\phi$ into discrete intervals of $\frac{2\pi}{50}=7.2^{\circ}$, while the elevation $\theta$ and the distance $z$ remain fixed. For each 3D instance, we replicate the same latent variable $z_{T}$ over 50 frames, which can minimize temporal flickering across different views. Additionally, we also set the parameter $\sigma_{t}$ in Eq. A.2 to zero, thereby further mitigating uncertainty introduced by varying noise patterns.

### v2

@@ 3.2 View Consistent Rendering
Given a source image $x^{\text{src}}$, our goal is to render a series of consistent novel views $x^{i}$. While the camera encoding technique of Sec. 3.1 significantly enhances pose accuracy, if images are sampled independently, they will almost never be visually consistent due to the intrinsic ambiguity of the reconstruction task. To remove or at least greatly mitigate this problem, we propose to sample images jointly.
@@
We first adapt the frame attention, a well-established method in video generators [58, 25, 7, 21, 72, 6], to capture temporal dependencies across views. As shown in Fig. 2 (c), given a 5D latent $z\in\mathbb{R}^{B\times v\times c\times h\times w}$, we initially reshape it to $z\in\mathbb{R}^{(B\times h\times w)\times v\times c}$, resulting in $batch\times height\times width$ sequences at the length of $views$. Subsequently, this reshaped latent is passed through the pseudo-3D attention module to calculate the similarity across different views. Since this attention layer operates across views but separately for each spatial location, the computational and memory costs are quite low (Tab. 1). Similar to ray conditioning, we inject multi-view attention at each level of the UNet $\hat{\epsilon}_{\theta}$.
@@
In order to reduce the variance between different views, we propose to start sampling each view from the same noise vector $x_{T}$. The network $\epsilon_{\theta}$ still generates different views because it is conditioned on the camera parameters. It can also generate different reconstructions by taking a new noise sample Fig. 7. However, sharing noise reduces the variance between views. This can be justified by noting that the network $\hat{\epsilon}_{\theta}(z_{t},t,y)$ is a continuous function of both $z_{t}$ and $y$ [71]. Besides, different from [66], we do not specially set the noise schedule in different time steps $t$. More complex designs and training strategies have the potential to improve performance but are not the focus of this work.
@@ 3.3 Learning formulation
where each training sample $(Z_{0},z^{\text{src}},\mathbf{P})$ consist of $N$ encoded target views $Z_{0}=\{\mathcal{E}(x^{i})\}_{i=1}^{N}$, the encoded source view $z^{\text{src}}=\mathcal{E}(x^{\text{src}})$ and the viewpoints $\mathbf{P}$. The network is conditioned on the source view and cameras and estimates the noise for all target views jointly. Following [38], we also concatenate the input image code $z^{\text{src}}$ to $z_{t}$ along the channel dimension and add the CLIP [50] encoding.
Figure 3: Perceptual Path Length Consistency (PPLC). To partly compensate for the viewpoint change, the second image is rectified w.r.t. the first before comparison. To illustrate the importance of using rectification, the figure shows two objects in a large azimuth $\phi:57.6^{\circ}$. The top row shows to the left an ideally-rendered image pair, which however attains a large LPIPS score due to the view change. To the right, rectification reduces this score. The bottom row shows the opposite, where a pair of incorrectly rendered views has its LPIPS increased by rectification.
@@ 3.4 Perceptual Path Length Consistency
To quantify the consistency across different views, recent methods [68, 38, 69, 76, 67] trained 3D models from the sampled views and evaluated them on the remaining views. However, this requires training a 3D model for each test instance, which is unfeasible for large-scale testing. Here, we propose to use instead the pairwise perceptual distance [80] to measure the consistency between generated views. In particular, we first subdivide the 360 rendering path into linear segments and then calculate the LPIPS score between two neighboring generated images $x^{i}$ and $x^{i+1}$. Naturally, these images differ due to the different viewpoints (Fig. 3). We partly compensate for the viewpoint change by rectifying [22] the second image w.r.t. the first. We then define Perceptual Path Length Consistency (PPLC) of a rendered sequence $\{x^{i}\}_{i=1}^{N}$ as follows:
@@ 4 Experiments
Table 1: Comparison with SoTA methods on all 7,729 objects in Objaverse test-set. Recent works like [38, 14, 69] were originally evaluated using a subset of this data only due to the cost of training additional 3D models. Unlike them, we directly evaluate the model on whole test-set without using additional 3D networks. The inference time is for rendering a single target view and a $360^{\circ}$ video, respectively.
@@ 4.1 Experimental Details
We follow [38, 69, 39] and assess the NVS quality by comparing the generated images and the ground-truth views at different levels of granularity, including PSNR, SSIM, LPIPS [80] [^1] https://github.com/richzhang/PerceptualSimilaritysqueeze”-net., and Fréchet Inception Distance (FID) [23] [^2] https://github.com/GaParmar/clean-fidclip_vit_b_32”-net.. Some of these metrics look at the statistics of individually reconstructed images rather than exact reconstructions, which is the correct approach given that the NVS task is inherently ambiguous; however, for the same reason, they are not suitable for assessing multi-view consistency. Instead, we use PPLC (Sec. 3.4) to evaluate consistency. To do so, we generate 50-frame videos along a pre-defined circular camera trajectory for each object and then compute the PPLC score between neighbouring frames.
@@
We compare Free3D to the state-of-the-art Zero-1-to-3 [38] and the follow-ups Zero123-XL [14] SyncDreamer [39] and Consistent123 [69]. While noticed other concurrent works in arXiv for NVS [77, 30, 76], but, as no codes are available during the submission, we do not re-train for comparison.
@@ 4.2 Assessing Quality
Table 3: Ablations of Free3D design choices. Here, to reduce the computational cost and testing time, we only evaluate 2,000 instances in Objaverse dataset, instead of running the whole dataset with 7,729 for 50-frame $360^{\circ}$ video rendering. Therefore, the PPLC score is slightly different from the values reported in Tab. 1.
@@
Qualitative results are visualized in Fig. 4. Free3D achieves better results even under challenging viewpoints. SyncDreamer [39] uses an explicit volumetric representation representation, but can only generate views with fixed elevation ($30^{\circ}$), leading to worse results on synthesizing arbitrary target views. While Consistent123 [69] aims to improve rendering consistency with a version of multi-view attention, they cannot directly improve pose accuracy. The Zero123-XL [14] learns to capture pose better than the baseline [38] by training on the larger Objaverse-XL, but the pose is still not very accurate. Because of the RCN layer, Free3D shows no such pose errors and results in better images.
@@ 4.3 Assessing Generalization
In Tab. 2, Free3D, which uses the same training set as Zero-1-to-3 [38], outperforms the baseline and all state-of-the-art variants, including Zero123-XL [14], which is trained on a larger 3D dataset, and SyncDreamer [39], which utilizes a heavier 3D volumetric representation. SyncDreamer improves the baseline on OmniObject3D, which has small elevation views change, while achieving worse results on GSO with random elevation angles. Consistent123 [69] is also trained multiple views jointly. However, their relative improvement is limited. Table 2 shows that Free3D achieves very substantial improvements on all instantiations on both OmniObject3D and GSO datasets. This further indicates that ray conditioning can successfully improve the pose accuracy of the target view.
@@
A qualitative comparison is given in Fig. 5 (more in supplemental Figs. C.3 and C.4). Though Free3D is only trained on objaverse dataset, it works quite well in the open-set setting. Moreover, it shows significantly better results than all state-of-the-art models.
Results on real case.
To further demonstrate the generalization of our Free3D to real images, following [36], we run Segment Anything [35] to extract the segments and show qualitative results in Fig. 1. Besides, we also show diverse results in Fig. 7 by sampling different noises.
Refer to captionReal ImageGenerated Views by our Free3D
Sample 1
Sample 2
Figure 7: Diverse NVS in real scenes. Zoom in to see the details.
@@ 4.4 Assessing 3D consistency
To assess 3D consistency, we render a $360^{\circ}$ video with the fixed elevation angle $\theta:=0^{\circ}$ and $50$ azimuth angles uniformly sampled in $[0^{\circ},360^{\circ}]$. The quantitative results are reported in Tabs. 1 and 2 using the proposed PPLC score, and the qualitative results are shown in Figs. 1 and 6. Consistent123 [69] still requires training additional NeRF to achieve 3D consistent video, while the directly rendered videos are still flickering. SyncDreamer [39] obviously improves view consistency by using an explicit 3D volume representation. However, it is trained only on a fixed elevation angle and can generate only 16 fixed frames for a video in its code. Interestingly, Free3D, by using multi-view attention and multi-view noise sharing, is nearly as effective (while being much cheaper). This is also partly due to ray conditioning, which can capture more precise poses of the target view and thus reduce ambiguity. The additional results are provided as videos in the supplemental materials.
@@ 4.5 Ablation Study
We run a number of ablations to analyse Free3D. Results are shown in Tab. 3 and discussed in detail next.
@@ Acknowledgements.
We use the Objaverse [15], OmniObject3D [73], and GSO [17] following their terms and conditions. These datasets contain synthetic or scanned 3D objects, but, as far as we could determine, no personal data. For further details on ethics, data protection, and copyright please see https://www.robots.ox.ac.uk/~vedaldi/research/union/ethics.html.
@@ Appendix B Experiment Details
360 Video Rendering.
To render a 360 video, we establish a circle trajectory by uniformly subdividing the azimuth $\phi$ into discrete intervals of $\frac{2\pi}{50}=7.2^{\circ}$, while the elevation $\theta$ and the distance $z$ remain fixed. For each 3D instance, we replicate the same latent variable $z_{T}$ over 50 frames, which can minimize temporal flickering across different views. Additionally, we also set the parameter $\sigma_{t}$ in Eq. A.2 to zero, thereby further mitigating uncertainty introduced by varying noise patterns.